OpenStudy (anonymous):

who can come up with the answer to: 1/5(9t-8)=11/5t- t+8/10 ive tried it several times and get the wrong answer...

zepdrix (zepdrix):

Hmm it's pretty difficult to read, lemme see if I can format it correctly.\[\large \frac{1}{5(9t-8)}=\frac{11}{5t}-\frac{t+8}{10}\]Is that what it's suppose to look like? I might be reading the last part incorrectly :O Let me know.

OpenStudy (anonymous):

yes, that is correct

OpenStudy (anonymous):

woops... nope, here it comes

OpenStudy (anonymous):